SURPRISING COMPLEXITY OF A SMALL MOLECULE: PHOTOFRAGMENTATION DYNAMICS OF ICN, ICN·Ar_{n} and ICN·(CO_{2})_{n}

\underline{ANNE B. McCOY}, Department of Chemistry, The Ohio State University, Columbus, OH 43210; AMANDA S CASE, JOSHUA P. MARTIN, and W. CARL LINEBERGER, JILA and Department of Chemistry and Biochemistry University of Colorado, Boulder, CO 80309.

The photofragmentation dynamics of ICN^{−} in isolation and in complexes with argon or CO_{2} are investigated through a combination of experimental and theoretical studies. Experimentally, we probe the excited state dynamics of ICN^{−} following excitation to the lowest energy state that correlates to I^{−} + CN products. In the absence of solvating atoms or molecules, most of the excess charge localizes on I^{−}, although ≈ 3% of the products have the charge localized on the CN^{−}. The introduction of a single CO_{2} or argon atom changes the branching ratio between the three possible product channels (I^{−} , CN^{−} or ICN^{−}). Interestingly when one argon atom is introduced roughly 5% of the products undergo cage recombination to form ICN^{−}, providing an example of single atom caging.^{a} When CO_{2} is introduced, cage recombination is only seen in larger complexes, but a significant fraction of the photoproducts obtained following excitation of ICN^{−}·CO_{2} have the charge localized on the CN. These results will be discussed and analyzed using potential surfaces evaluated at the MR-SOCISD level of theory.^{b}

aA. S. Case, E. M. Miller, J. P. Martin, Y.-J. Lu, L. Sheps, A. B. McCoy and W. C. Lineberger, Angew. Chem. Int. Ed. 51, 2651 (2012)

bA. B. McCoy, IJQC, (in press).